# Implementation of an ITER-relevant QP-based Current Limit Avoidance algorithm in the TCV tokamak

### v1

@@ 1 Introduction
The tokamak concept is still considered a baseline for designing a potential power plant reactor based on thermonuclear fusion [1]. The next step in this direction is the International Thermonuclear Experimental Reactor (ITER) [2], which will extensively test concepts and technologies required for the design and construction of a fusion reactor. A tokamak is a toroidal-shaped device in which a fully ionized gas of hydrogen isotopes is heated up to temperatures of hundreds of millions of degrees Celsius, where the thermal agitation of the electrically charged nuclei can overcome the Coulomb repulsive force. This makes possible collisions producing fusion reactions. External magnetic fields allow the confinement of such a hot plasma. These fields are generated using currents flowing in the Poloidal Field (PF) and the Toroidal Field (TF) coils. These are shown in a simplified schematic for the TCV tokamak in Fig. 1.
@@
An important role in tokamaks is played by magnetic control which is needed since day one and is essential to successfully control high-performance plasmas, such as those envisaged for ITER [3].
@@
As a consequence, it is important for a reliable plasma magnetic control system to include a Current Limit Avoidance (CLA) system to safely operate the machine also when disturbances and/or uncertainties drive the PF currents close to their limits [15, 16, 17]. In fact, for a critical infrastructure such as a nuclear power plant, common-mode failures must be avoided by adopting different hardware and software solutions. In [18], three different limit avoidance algorithms are described.
@@
Other possible approaches to limit currents in the active circuits resemble Model Predictive Control (MPC, [19, 7]). However, while the implementation of a MPC scheme typically needs the knowledge of the plant dynamics, this is not the case for the CLA approach, provided that a so-called current control scheme is adopted [5]. In addition, optimization problems arising in MPC control schemes are usually larger than those solved in the CLA formulation proposed in this article, as they must take into account more than one prediction step. This makes the CLA architecture less demanding from a computational point of view.
@@
The CLA strategy has proven its effectiveness in numerical simulations on the DEMO device [18, 20], and has been proposed for inclusion in the ITER Plasma Control System (PCS) [21]. In this paper, the current limit avoidance system proposed for ITER is tested as a proof of concept on TCV, a medium-sized tokamak operated by the Swiss Plasma Center of the École Polytechnique Fédérale de Lausanne. The flexibility of TCV’s upgraded Distributed Control System (Système de Contrôle Distribué, or SCD [22]) enables rapid and effective implementation and testing of novel plasma control solutions, including a recently implemented shape control algorithm [23, 24] and an alternative version of the CLA system which relies on a combination of a dynamic optimizer and a static annihilator [25]. The significantly faster dynamics observed in TCV plasmas with respect to the ones expected in larger devices such as ITER require higher sampling frequencies in the control system, making the real-time implementation more challenging; this in turn implies that solutions adopted on TCV are expected to be easily scalable to future reactors, at least from the computational point of view, making TCV an ideal testbed for tokamak control technologies. One of the main differences between ITER and TCV is in the technology of the coils. TCV is not a superconductive tokamak, which means that resistive copper coils are used to confine the plasma, while for ITER a set of superconductive coils is foreseen. However, the proposed CLA system only uses steady-state information, allowing for a proper validation of the proposed methodology despite the difference in the coil dynamics between the TCV and ITER devices.
@@ 2 TCV magnetic control system
A layout of the TCV poloidal cross-section is shown in Fig. 2. The core magnetic controller in the TCV digital control system is designed to emulate the legacy analog system, called the hybrid controller. The name comes from the fact that the original system was a hybrid digital-analog controller, where the computations were carried out analogically, but the gain matrices could be switched digitally. The hybrid controller, represented as the orange block in Fig. 3, is a Multiple-Input Multiple-Output (MIMO) controller that takes care of the vertical stabilization of the plasma and of the regulation of the plasma current $I_{p}$, of the radial and vertical plasma centroid displacement, and of tracking scenario coil currents. In particular, the average current in the $OH$ coils and two combinations of the $EF$ ones are dedicated to $I_{p}$ control and to plasma position control, respectively. The remaining currents are projected in a space that is orthogonal to these combinations; for details, see [24]). The hybrid controller generates voltage requests for the PF coils, which are summed to feedforward traces $V_{ff}$, computed offline, to obtain the voltage requests $V_{a}$ to the coil power supplies.
Notice that, to be able to stabilize the plasma vertically, the digital version of this controller needs to run at a rather high sampling rate, typically chosen equal to $10$ kHz.
As it is common in many tokamaks, the controller relies on magnetic and current measurements to reconstruct plasma position and current. In addition, a fast observer of the vertical position is obtained by using the in-vessel magnetic probes only, which are less affected by the shielding effect of the passive structures, and this is used by the vertical stabilization control loop.
@@
The hybrid controller also includes a channel to control the plasma density, reconstructed through interferometry, which in practice is completely decoupled from the magnetic controller. A detailed analysis of the hybrid controller is given in [24]. An emulator of this controller is available in Matlab-Simulink, allowing to simulate the behavior of the controller before its real-time application. Moreover, the framework allows for automatic code generation and deployment of control algorithms directly in the SCD system [26], providing a friendly environment to extend the capabilities of the original controller.
@@
Furthermore, the TCV magnetic control system is equipped with a plasma shape controller, as shown in Fig. 3; the design and validation of such controller are extensively discussed in [23, 24]. Unlike the hybrid controller, the feedback signals for the shape controller, which is based on the isoflux approach and modifies the reference signals of the hybrid block, are obtained by interpolating the poloidal flux and magnetic field maps obtained from the real-time equilibrium reconstruction code LIUQE [27]. In practice, the necessity of a full equilibrium reconstruction limits the sampling rate of the shape controller to $1$ kHz.
@@
Finally, note that the coils used for plasma shaping in TCV are the sets E1-8 and F1-8 shown in Fig. 2 in red and blue, respectively. Therefore, the associated currents $I_{EF}$ are the ones that will be considered in the allocation problem solved by the CLA.
@@ 3 The Current Limit Avoidance Algorithm
Figure 3: TCV control scheme: in orange the hybrid magnetic controller, including a control loop for the PF currents; in green the shape controller; in blue the proposed CLA system.
@@
The proposed CLA system, shown in blue in Fig. 3, acts in parallel to the shape controller, providing additional control actions, summed to the hybrid controller reference signals $r_{h}$ with the aim of optimizing a quadratic cost function while simultaneously satisfying a set of hard linear constraints. Moreover, the steady-state effect of the allocator on the plasma shape is subtracted from the reference signals $r_{sh}$, to avoid any conflict between the two systems. The objective of the proposed CLA system is to reduce any excess of currents with respect to saturation limits by resorting to an alternative control pattern that allows the re-allocation of the currents while minimizing the degradation in terms of the shape control performance.
@@
The CLA receives as input the references $r_{h}$ to the hybrid controller, modified by the action of the shape controller $u_{sh}$. As discussed in Section 1, the proposed allocator acts based on the steady-state effect of the shape control inputs on the coil currents. For this reason, the first step is to obtain an estimate of their steady-state value. To do so, the proposed CLA system takes into account the static gain of the closed-loop system consisting of the linearized TCV plasma response in the vicinity of the considered plasma configuration and a state-space model of the hybrid controller. Let us denote the static gain from (variations of) the hybrid control error $e_{h}$ to the currents in the E and F coils of TCV as $P^{0}_{EF}$ and the gain from $e_{h}$ to variations in the plasma shape descriptors $y_{sh}$ as $P^{0}_{sh}$. Then, assuming that the plasma is close to the desired scenario configuration and denoting with $y_{h}$ the outputs of the plant controlled by the hybrid loop, an estimate of the steady-state value of the currents in the E and F coils of TCV can be obtained at each controller step as
@@
$I^{0}_{EF}=\underbrace{I_{EF}}_{\begin{subarray}{c}\text{measured}\\ \text{currents}\end{subarray}}+\,\,\underbrace{P^{0}_{EF}(u_{sh}+r_{h}-y_{h})}_{\begin{subarray}{c}\text{expected variation}\\ \text{at steady-state}\end{subarray}}\,.$ (1)
Similarly, the steady-state effect of the CLA on the plasma shape descriptors $y_{sh}$ can be estimated as
$\delta y^{0}_{sh}=P^{0}_{sh}u_{CLA}\,,$ (2)
while the steady-state effect of the CLA only on the $I_{EF}$ currents is
$\delta I^{0}_{EF}=P^{0}_{EF}u_{CLA}\,.$ (3)
@@
$\min_{u_{CLA}}\qquad\delta y^{0T}_{sh}W\delta y^{0}_{sh}+\delta I^{0T}_{EF}Q\delta I^{0}_{EF}\,,$ (4)
@@
where $\underline{I}_{EF}$ and $\overline{I}_{EF}$ represent the lower and upper bounds on the $EF$ coil currents, respectively. The first term of the cost function aims at minimizing the effect of the CLA on the shape descriptors $y_{sh}$, while the second aims at reducing its effect on the PF coil currents. Hence, if the steady-state current vector $I^{0}_{EF}$ satisfies the constraints of the optimization problem, the optimal solution is given by $u_{CLA}~{}=~{}0$, i.e., the CLA leaves the references to the hybrid controller $r_{h}$ unmodified.  [^1] Alternatively, the operator may decide to minimize the $I_{EF}$ steady-state current norm by changing the second term in the optimization to $I^{0T}_{EF}QI^{0}_{EF}$, similarly to what is proposed in [25]. In this case, typically a relatively low weight $Q$ is assigned to the current term with respect to the weight $W$ assigned to the shape variations, and the optimizer will attempt to rearrange the currents in order to obtain a comparable shape accuracy while minimizing the weighted current norm as a secondary objective.
@@
An important remark is that in the case of TCV and differently to what has been discussed in Section 1, the design of the CLA still (loosely) depends on the knowledge of the structure of the hybrid controller, and in particular of its steady-state gain. This is due to the fact that, in TCV, the control of the coil currents and the plasma current and position is tightly coupled, while the shape control also generates reference variations for the plasma position channels. On the other hand, in other tokamak magnetic control architectures, where the plasma position and $I_{p}$ controllers generate references for an internal loop that regulates the currents in the PF coils, this knowledge is not necessary under the assumption that the coil current controller has an approximately unitary steady-state gain on all channels, i.e., that the steady-state values of the PF currents are close to the respective references. This is, for example, the case of the foreseen ITER control system.
Finally, notice that, in other formulations of the same problem (such as [18]), the effect of the CLA on the plasma current is also taken into account explicitly through an additional term in the optimization cost function. However, in TCV the $I_{p}$ control is achieved by a dedicated set of coils (OH1-2 in Fig. 2) on time scales that are significantly faster than those of the shape control, so this term can be neglected in the case under exam.

### v2

@@ 1 Introduction
The tokamak concept is still considered a baseline for the design of a potential power plant reactor based on thermonuclear fusion [1]. The next step in this direction is ITER [2], which will extensively test concepts and technologies required for the design and construction of a fusion reactor. A tokamak is a toroidal-shaped device in which a fully ionized gas of hydrogen isotopes is heated up to temperatures of hundreds of millions of degrees Celsius, where the thermal agitation of the electrically charged nuclei can overcome the Coulomb repulsive force. This enables collisions producing fusion reactions. External magnetic fields allow the confinement of such a hot plasma. These fields are generated using currents flowing in the Poloidal Field (PF) and the Toroidal Field (TF) coils. These are shown in a simplified schematic for the TCV tokamak in Fig. 1. An important role in tokamaks is played by magnetic control which is needed since day one and is essential to successfully control high-performance plasmas, such as those envisaged for ITER [3].
@@
As a consequence, it is important for a reliable plasma magnetic control system to include a Current Limit Avoidance (CLA) system to safely operate the machine when disturbances and/or uncertainties drive the PF currents close to their limits [15, 16, 17]. In fact, for a critical infrastructure such as a nuclear power plant, common-mode failures, i.e. failures that occur simultaneously in several components or subsystems due to a shared cause, must be avoided. This can be done by adopting different alternative hardware and software solutions to improve the overall level of reliability of the system. To this end, in [18], three different limit avoidance algorithms are described and tested through numerical simulations for the DEMO tokamak. Two of these algorithms solve an online optimization problem, based on Quadratic Programming (QP) or Linear Programming (LP) respectively, while the third exploits a dynamic optimizer similar to what is proposed in [16, 19]. These systems allow to modify the PF reference currents to satisfy the saturation limits while reducing the impact on the plasma shape by exploiting actuator redundancy. In this article, we aim at demonstrating the experimental feasibility of one of such solutions. In particular, a variant of the QP solution proposed in [18] is implemented and experimentally validated on TCV.
@@
Another possible approach which is often considered to limit currents in the active circuits is Model Predictive Control (MPC, [20, 7]). However, while the implementation of a MPC scheme typically needs the knowledge of the plant dynamics, this is not the case for the CLA approach, provided that a so-called current control scheme is adopted [5]. In addition, optimization problems arising in MPC control schemes are usually larger than those solved in the CLA formulation proposed in this article, as they take into account more than one prediction step. This makes the CLA architecture less demanding from a computational point of view.
@@
The CLA strategy has proven its effectiveness in numerical simulations on the DEMO device [18, 21], and has been proposed for inclusion in the ITER Plasma Control System (PCS) [22]. In this paper, the current limit avoidance system proposed for ITER is tested as a proof of concept on TCV, a medium-sized tokamak operated by the Swiss Plasma Center of the École Polytechnique Fédérale de Lausanne. The flexibility of TCV’s upgraded Distributed Control System (Système de Contrôle Distribué, or SCD [23]) enables rapid and effective implementation and testing of novel plasma control solutions, including a recently implemented shape control algorithm [24, 25] and an alternative version of the CLA system which relies on a combination of a dynamic optimizer and a static annihilator [19]. The significantly faster dynamics observed in TCV plasmas with respect to the ones expected in larger devices such as ITER require higher sampling frequencies in the control system, making the real-time implementation of complex control algorithms more challenging. This in turn implies that solutions adopted on TCV are expected to be easily scalable to future reactors, at least from the computational point of view, making TCV an ideal testbed for tokamak control technologies. One of the main differences between ITER and TCV is in the technology of the coils. TCV is not a superconductive tokamak, which means that resistive copper coils are used to confine the plasma, while for ITER a set of superconductive coils is foreseen. However, the proposed CLA system only uses steady-state information, allowing for a proper validation of the proposed methodology despite the difference in the coil dynamics between the TCV and ITER devices. Moreover, in TCV not all of the PF coil currents are controlled explicitly, but instead some combinations of currents are dedicated to plasma position and current control. This requires some adjustments in the QP-CLA strategy proposed in [18], which are discussed in section 3.
@@ 2 TCV magnetic control system
A layout of the TCV poloidal cross-section is shown in Fig. 2. The core magnetic controller in the TCV digital control system is designed to emulate the legacy analog system, called the hybrid controller. The name comes from the fact that the original system was a hybrid digital-analog controller, where the computations were carried out analogically, but the gain matrices could be switched digitally. The hybrid controller, represented in Fig. 3, is a Multiple-Input Multiple-Output (MIMO) controller that takes care of the vertical stabilization of the plasma and of the regulation of the plasma current $I_{p}$, of the radial and vertical plasma centroid displacement, and of tracking scenario coil currents. In particular, the average current in the OH coils and two combinations of the E and F ones are dedicated to $I_{p}$ control and to plasma position control, respectively. The remaining currents are projected in a space that is orthogonal to these combinations; in particular, the difference $I_{DOH}=I_{OH,1}-I_{OH,2}$ is usually controlled to zero to minimize the stray field, while $14$ independent combinations of E and F currents, denoted by $I_{orth}$, are controlled explicitly (for more details, see [25]). The hybrid controller generates voltage requests for the PF coils, which are summed to feedforward traces $V_{ff}$, computed offline, to obtain the voltage requests $V_{a}$ to the coil power supplies.
@@
As is common in many tokamaks, the controller relies on magnetic and current measurements to reconstruct plasma position and current. In addition, a fast observer of the vertical position is obtained by using the in-vessel magnetic probes only, which are less affected by the shielding effect of the passive structures; this observer is used by the vertical stabilization control loop. Notice that, to be able to stabilize the plasma vertically, the digital version of this controller needs to run at a rather high sampling rate, typically chosen equal to $10$ kHz. The hybrid controller also includes a channel to control the plasma density, reconstructed through interferometry, which in practice is completely decoupled from the magnetic controller. An emulator of this controller is available in Matlab-Simulink, allowing to simulate the behavior of the controller before its real-time application. Moreover, the framework allows for automatic code generation and deployment of control algorithms directly in the SCD system [26], providing a user-friendly environment to extend the capabilities of the original controller.
@@
Furthermore, the TCV magnetic control system is equipped with a plasma shape controller, as shown in Fig. 3; the design and validation of such controller are extensively discussed in [24, 25]. The considered shape controller is based on the isoflux approach [27]. To achieve the desired shape, a set of flux differences between control points specified by the user and a chosen boundary-defining point, such as an X-point or a limiter point, are controlled to zero. Moreover, the field components at the desired X-points are also controlled to zero in the case of diverted plasma configurations, while for limiter plasmas the component of the magnetic field tangent to the wall at the limiter point can be controlled to zero. The feedback signals for the shape controller, which modifies the reference signals of the hybrid block, are obtained by interpolating the poloidal flux and magnetic field maps obtained from the real-time equilibrium reconstruction code LIUQE [28]. In practice, the necessity of a full equilibrium reconstruction limits the sampling rate of the shape controller to $1$ kHz.
@@
Finally, note that the coils used for plasma shaping in TCV are the sets E1-8 and F1-8 shown in Fig. 2 in green and blue, respectively. Therefore, the associated currents $I_{EF}$ are the ones that will be considered in the allocation problem solved by the CLA.
@@ 3 The Current Limit Avoidance Algorithm
Figure 3: TCV control scheme, including the hybrid magnetic controller, the shape controller and the proposed CLA system (in light blue).
@@
The proposed CLA system, shown in light blue in Fig. 3, acts in parallel to the shape controller, providing additional control actions, summed to the hybrid controller reference signals $r_{h}$ with the aim of optimizing a quadratic cost function while simultaneously satisfying a set of hard linear constraints. Moreover, the steady-state effect of the allocator on the plasma shape is subtracted from the reference signals $r_{sh}$, to avoid any conflict between the two systems. The objective of the proposed CLA system is to reduce any excess of currents with respect to saturation limits by resorting to an alternative control pattern that allows the re-allocation of the currents while minimizing the degradation in terms of the shape control performance.
@@
The CLA receives as input the references $r_{h}$ to the hybrid controller, modified by the action of the shape controller $u_{sh}$. As mentioned in section 1, the proposed allocator acts based on the expected steady-state effect of the shape control actions on the coil currents. For this reason, the first step is to obtain an estimate of this relation. Since, as explained in section 2, in TCV some coil current directions are not explicitly regulated, but are rather directly allotted to the control of the plasma position and current, to obtain such an estimate the proposed CLA system takes into account the static gain of the closed-loop system consisting of the linearized TCV plasma response in the vicinity of the considered plasma configuration and a state-space model of the hybrid controller. An estimate of said static gain can be computed based on linearized models obtained through the linearization module of the fge code [29], included in the MEQ suite, similarly to what is done in [24, 25, 19]. In this way, variations in the references to the plasma position channels can be translated into expected variations in the coil currents. It is worth to notice that, for the purposes of this article, the control channels associated to $I_{p}$ and to the current in the OH coils are considered as decoupled from those taking care of the plasma position and of the E and F coils, and as such they are not considered in the proposed architecture.
Let us denote the static gain from (variations of) the hybrid control errors $e_{h}$ to the currents in the E and F coils of TCV as $P^{0}_{EF}$ and the gain from $e_{h}$ to variations in the plasma shape descriptors $y_{sh}$ as $P^{0}_{sh}$. The vector $e_{h}$ includes control errors on the $I_{orth}$ combinations of the E and F coil currents introduced in section 2 and on the plasma centroid position. The shape descriptors $y_{sh}$, instead, include the magnetic field components at the desired X-points and the differences between poloidal flux at a set of user-specified contour points and a boundary-defining point (e.g. the X-point for the diverted configuration considered in the examples), according to the well-established isoflux control paradigm.
Assuming that the plasma is close to the desired scenario configuration and denoting with $y_{h}$ the outputs of the plant controlled by the hybrid loop, an estimate of the steady-state value of the currents in the E and F coils of TCV can be obtained at each controller step as
@@
$I^{0}_{EF}=\underbrace{I_{EF}}_{\begin{subarray}{c}\text{measured}\\ \text{currents}\end{subarray}}+\,\,\underbrace{P^{0}_{EF}(u_{sh}+{u_{CLA}}+r_{h}-y_{h})}_{\begin{subarray}{c}\text{expected variation}\\ \text{at steady-state}\end{subarray}}\,,$ (1)
For simplicity, with reference to Fig. 3, in what follows we will assume $V_{ff}=const$ during the shape control phase. This is not the case in standard TCV shots, but the voltage feedforwards were frozen to their value at the controller activation times for the experiments reported in this article. Note that the introduction of the static gain matrix $P_{EF}^{0}$ is an approximation used to adapt the logic proposed in [18] to TCV, where the control channels for the PF coil currents and for the plasma position are not independent. To this end, the proposed adaptation relies on the assumption that the hybrid controller is accurate enough to bring the controlled errors close to zero at steady-state. In practice, this means that, after the transient phase, $I_{EF}^{0}\approx\lim_{t\to\infty}I_{EF}$ in eq. (1). Note that if the hybrid controller is capable of accurate reference tracking and if the plasma configuration is not significantly perturbed by the action of the allocator, then the term in round brackets in (1) is expected to be small and the linearized response approximation holds.
It is worth remarking however that the proposed architecture in practice only acts on the references to the inner control loops, and hence it is only accurate to the extent to which the underlying control loop (the hybrid controller in this case) is. In fact, if a steady-state error was present in the quantities tracked by such a controller, this would be reflected in the steady-state currents achieved, i.e. $I_{EF}^{0}\neq\lim_{t\to\infty}I_{EF}$. More details regarding this point are discussed in section 4.2.
On the other hand, it is also worth to note that, if an explicit control of the PF coil currents is used (as foreseen in ITER’s magnetic control architecture), the proposed allocator can directly act on the references to such controller, and hence the approximate estimate of the steady-state currents given by (1) is not necessary. This is what is done, for example, in the approach proposed in [18], where the CLA acts on the current references fed to an inner control loop for the coil currents. As it happens with the solution proposed in the present work, the allocator in [18] is also agnostic with respect to the steady-state error of the coil current controller. In practice, the introduction of the static gain $P_{EF}^{0}$ in 1 is a device used to adapt the logic proposed for ITER in [18] to TCV, where the control channels for the PF coil currents and for the plasma position are not independent.
We are interested in the variations of the shape descriptors and the EF currents induced by the action of the proposed allocation strategy once a steady-state has been reached. These are denoted as $\delta y^{0}_{sh}$ and $\delta I^{0}_{EF}$, respectively. The steady-state effect of the CLA on the plasma shape descriptors $y_{sh}$ and on the EF coil currents can be estimated as
$\delta y^{0}_{sh}=P^{0}_{sh}u_{CLA}\,,\quad\delta I^{0}_{EF}=P^{0}_{EF}u_{CLA}\,.$ (2)
@@
$\min_{u_{CLA}}\qquad$ $\delta y^{0T}_{sh}W\delta y^{0}_{sh}+\delta I^{0T}_{EF}Q\delta I^{0}_{EF}$ (3)
s.t. $\underline{I}_{EF}\leq I^{0}_{EF}\leq\overline{I}_{EF}$
where $\underline{I}_{EF}$ and $\overline{I}_{EF}$ represent the lower and upper bounds on the $EF$ coil currents, respectively. The first term of the cost function aims at minimizing the effect of the CLA on the shape descriptors $y_{sh}$, while the second aims at reducing its effect on the PF coil currents. The weights $W$ and $Q$ are properly scaled to consider the magnitude difference between shape and active currents values. Note that, if the steady-state current vector $I^{0}_{EF}$ satisfies the constraints of the optimization problem, the optimal solution is given by $u_{CLA}~{}=~{}0$, i.e. the CLA leaves the references to the hybrid controller $r_{h}$ unmodified.  [^1] On the other hand, the operator may decide to minimize the $I_{EF}$ steady-state current norm by changing the second term in the optimization to $I^{0T}_{EF}QI^{0}_{EF}$, similarly to what is proposed in [19]. In this case, typically a relatively low weight $Q$ is assigned to the current term with respect to the weight $W$ assigned to the shape variations, and the optimizer will attempt to rearrange the currents in order to obtain a comparable shape accuracy while minimizing the weighted current norm as a secondary objective.
Substituting equations (1) and (2) into (3) we obtain the following standard QP problem with linear constraints
$\min_{u_{CLA}}\qquad$ $(P^{0}_{sh}u_{CLA})^{T}W(P^{0}_{sh}u_{CLA})+(P^{0}_{EF}u_{CLA})^{T}Q(P^{0}_{EF}u_{CLA})$ (4)
@@
An important remark is that in the case of TCV and differently to what has been discussed in section 1, the design of the CLA still (loosely) depends on the knowledge of the structure of the hybrid controller, and in particular of its steady-state gain. This is due to the fact that, in TCV, the control of the coil currents and the plasma current and position is tightly coupled, while the shape control also generates reference variations for the plasma position channels. On the other hand, in other tokamak magnetic control architectures, where the plasma position and $I_{p}$ controllers generate references for an internal loop that regulates the currents in the PF coils, this knowledge is not necessary under the assumption that the coil current controller has an approximately unitary steady-state gain on all channels, i.e., that the steady-state values of the PF currents are close to the respective references. This is, for example, the case of the foreseen ITER control system. As a consequence, the implementation of the algorithm is also expected to be simpler, as the effect of the position controller on the PF coil currents must not be taken explicitly into account.
However, it is important to note that the performance of the proposed allocation strategy, which acts similarly to a reference governor and relies on the availability of a model of the relationship between the PF currents and the plasma shape, strongly depends on the accuracy of the available models and on the underlying control architecture. As the codes used to produce the results in this paper are available to the fusion community, there is no reason to expect lower modeling accuracy on ITER compared to TCV. Further details on the impact of the accuracy of the underlying controller on the performance of the proposed strategy are given instead in section 4.2.
Finally, it should be mentioned that, in other formulations of the same problem (such as [18]), the effect of the CLA on the plasma current is also taken into account explicitly through an additional term in the optimization cost function. However, in TCV the $I_{p}$ control is achieved by a dedicated set of coils (OH1-2 in Fig. 2) on time scales that are significantly faster than those of the shape control, so this term can be neglected in the case under exam.